My question is from a review sheet from last year not in our book. I need to understand how to find the surface area of a cylindrical can. If the can is 15cm high and with a diameter of 8cm. Thanks
Answer by checkley71(8403)

SA=2pieR~2+2pieRH OR SA=2*3.14*8*8+2*3.14*8*15 OR SA=401.92+376.8 OR SA=778.7CM~2.